The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
There is no y coordinate to go with (2,). This is a test question and this is how it was written by the teacher. Could you please help me!
: The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
There is no y coordinate to go with (2,). This is a test question and this is how it was written by the teacher. Could you please help me!

Answer by stanbon(19673)

The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
--------------
a. (8,6) is a point of the equation
b. that is not an equation
c. (8,6) is a point of the equation
d. (8,6) is not a point of the equation
---------------
Let x = 2 in "a" and in "c" you get the following:
In a: y = -4
In c: y = 21/2
==============================
Cheers,
Stan H.
Question 148090: The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
There is no y coordinate to go with (2,). This is a test question and this is how it was written by the teacher. Could you please help me!
: The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
There is no y coordinate to go with (2,). This is a test question and this is how it was written by the teacher. Could you please help me!

Answer by oscargut(682)

yes it is strange, options could be a) and c) but without y coordinate of (2,) you can not do it
Question 148090: The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
There is no y coordinate to go with (2,). This is a test question and this is how it was written by the teacher. Could you please help me!
: The equation of the line through (8,6) and (2,) is
a. 5x-3y=22
b. 3/5x + 8/5
c. 3x + 4y = 48
d. -4x + 2y = -16
There is no y coordinate to go with (2,). This is a test question and this is how it was written by the teacher. Could you please help me!

Answer by nerdybill(1279)

In this case, since they gave you a single point at (8,6), apply that point to each of the given equations:
a. 5x-3y=22
5(8)-3y=22
40-3y=22
-3y=-18
y = 6
.
b. 3/5x + 8/5
3/5(8) + 8/5
24/5 + 8/5
32/5 (Does NOT equal 6 -- toss it out)
.
c. 3x + 4y = 48
3(8) + 4y = 48
24 + 4y = 48
4y = 24
y = 6
.
d. -4x + 2y = -16
-4(8) + 2y = -16
-32 + 2y = -16
2y = 32
y = 16 (Does NOT equal 6 -- toss it out)
.
Right now, it could either be 'a' or 'c'
a. 5x-3y=22
c. 3x + 4y = 48
.
I would think, at this point, that you could possibly have a typo when you keyed it in. Can you verify that the problem is typed in correctly?
.
Or, if you proceeded, the other piece of information given was (2,)
we could try to plug it into both 'a' and 'c':
a. 5x-3y=22
5(2)-3y=22
10-3y=22
-3y=12
y = -4
.
c. 3x + 4y = 48
3(2) + 4y = 48
6 + 4y = 48
4y = 42
y = 10.5